Mattel, Inc. is expected to pay a $1.60 dividend per share annually. Estimate its intrinsic value per common share using the dividend discount model (DDM) under each of the following separate assumptions.

(Assume that Mattel’s cost of equity capital is 8.0%.)

Required

a. The $1.60 dividend per share occurs at the end of each of the next three years, after which there are no additional dividend payments.

Round answer to two decimal places.

$Answer


b. The $1.60 dividend per share occurs at the end of each year in perpetuity.

Round answers to two decimal places, if applicable.

$Answer


c. The $1.60 dividend per share occurs at the end of each of the next three years, after which the dividends increase at a rate of 4% per year.

Round answers to two decimal places.

$Answer
Answer #1

Answer:

Expected dividend per share = $1.60
Cost of capital = 8%

a) Value of share = 1.60/1.08 + 1.60/(1.08)2 + 1.60/(1.08)3
=1.48+1.37+1.27
=$4.12 per share

b) Value of share = 1.60/8% =$20 per share

c) value of share =

Year Dividend PVF @ 8% PV value
1 $1.60 0.926 1.48
2 $1.60 0.857 1.37
3 $1.60 0.794 1.27
Tv $52 0.794 41.29
$45.41

Therefore value per share = $45.41

*Terminal value = D4/ke-g =1.60 x 1.04 / 0.008-0.04 =$52